Assume the following information:
Spot rate of Canadian dollar : $.80
90-day forward rate of Canadian dollar : $.79
90-day Canadian interest rate : 4%
90-day U.S. interest rate : 2.5%
a) What would be the return to a U.S. investor who used covered interest arbitrage from investing
in Canada? (assume the investor invests $1,000,000). Does the return exceed the return from
investing in the U.S. over the 90-day period? Is it worthwhile for the U.S. investor to invest in
Canada?
b) Explain completely the market forces on the spot market and the forward market to eliminate
the covered interest arbitrage
Answer #1

a) 90-day U.S. interest rate : 2.5%

Invest US$1,000,000 in Canada

Spot rate of Canadian dollar: $.80

US$1,000,000 = US$1,000,000/0.80 = Canadian $1,250,000

Interest earned on Canadian $1,250,000 * 0.04 = C$50000

Return on Canadian investment = C$50000+C$1,250,000 = C$1300000

90-day forward rate of Canadian dollar, convert to US$ : $.79 = C$1300000*$.79 = $1,027,000

Total return on arbitrage = ($1,027,000-US$1,000,000) / (US$1,000,000) = 2.7%, the yield is greater than US interest rate, so investor can avail this arbitrage opportunity.

b. Any increase in US interest rate with decrease in canadian interest rate and any increase in Canadian dollar sport rate with decrease in its forward rate would eliminate the covered interest arbitrage opportunity benefits.